OpenStudy (anonymous):

Write the equation of a line that passes through the points (-5, 9) and (-4, 7). y = 2x - 19 y = 2x - 1 y = -2x + 19 y = -2x – 1

OpenStudy (anonymous):

to find slope you do (y2-y1)/(x2-x1) (9-7)/(-5+4) 2/-1 -2 to find equation y-y1=m(x-x1) y-9=-2(x+5) distribute y-9=-2x-10 add 9 y=-2x-1

OpenStudy (anonymous):

y = 2x -1.